Prince can consume 2 goods X¹ and X² at a price p¹ and p² respectively .Princes preferences are represented by the following utility function U¹(X¹,X²) =X¹ X² and has an income of M .


a) Derive Princes Marshallian demand functions for the 2 goods

b)What are the Marshallian demand curves for the 2 goods if Princes utility changes to U²(X¹ ,X²)=⅛(X¹ ,X²).

c)Derive an expression for the marginal rate of substitution (MRS) between X¹ and X² for U²(X¹ ,X²)=⅛(X¹ ,X²).

e) suppose a production function is given by f(k,l) =l²k ,the price of capital is $10 and of labour $15.What combination of labour and capital minimises the cost of producing any given output

f) The production function of a given product is given by q=50kl. If the price of capital is $120 per day and the price of labour is $30 per day

What is the minimum cost of producing 1000 units of output ?


1
Expert's answer
2020-07-15T12:58:44-0400

a) "U1'=X1+X2\ufeff"

b)



c) "U'(x1)=MU1=\frac{X2}{8}"

"U'(x2)=MU2=\frac{X1}{8}"

"MRS=\frac{MU1}{MU2}=\frac{X2}{X1}"

e) "M=k"

"k=I^2k"

"I=1"

f) 50-days, lk-cost of production in day,we need 1000 units

"1000=50*lk"

"20=lk"